# Exploratory study on the masses of odd-$Z$ nuclei and $r$-process simulation based on the deformed relativistic Hartree-Bogoliubov theory in continuum

### v1

@@ IV $R$-process simulation
The obtained pseudo DRHBc mass table in Section III is applied in the $r$-process simulation, to study the impact of deformation effects in the $r$-process. The $r$-process simulation is performed using the classical $r$-process model, where nuclear masses are taken from the pseudo DRHBc mass table or the RCHB mass table [25], if the experimental data [8] are not available. For the $\beta$-decay rates, the empirical formula [99] using decay energies from the pseudo DRHBc or RCHB calculations is employed together with experimental data [100]. The astrophysical trajectory, i.e., the weight $\omega$ and the irradiation time $\tau$ of neutron flows, is determined by fitting the obtained abundances to the solar $r$-process abundances [101] at a temperature of $T=1.5\ \mathrm{GK}$.
The $r$-process abundances with pseudo DRHBc and RCHB mass tables are shown in Fig. 4. It is found that the simulated abundances with both the RCHB and pseudo DRHBc mass tables exhibit abundance peaks near $A=130,195$ and an abundance trough around $A=170$. The $r$-process abundances based on pseudo DRHBc mass table are higher for $A=148-165$ and lower for $A=170-178$ compared to those based on the RCHB mass table. This indicates the impact from the deformation effects.
@@
Figure 4: The $r$-process abundances from classical $r$-process simulations based on pseudo DRHBc and RCHB mass tables as functions of the mass number $A$. The solar $r$-process abundances [101] are displayed with black dots.
@@
The most important nuclei for the $r$-process are located in the $r$-process path. The $r$-process paths for nuclei with $A\sim 130-195$ based on the pseudo DRHBc and RCHB mass tables are displayed in Fig. 5. It is found that the $r$-process path nuclei around $A=160$ predicted by the pseudo DRHBc mass table are closer to the stability line than the ones predicted by the RCHB mass table. Since the beta decay half-lives are generally shorter when the nuclei are moving toward the neutron-rich direction, the decay half-lives of $r$-process path nuclei around $A=160$ based on the pseudo DRHBc mass table are longer than those based on the RCHB mass table. This results in more nuclei accumulating on the path nuclei. It is therefore understandable that the $r$-process abundances based on the pseudo DRHBc mass table are larger for $A=148-165$ than the ones based on the RCHB mass table. On the other hand, it is noted that the $r$-process path predicted by the pseudo DRHBc mass table has a gap around $A=175$, which leads to a deficiency in the simulated abundances in this region.
@@
Figure 5: The $r$-process path for nuclei with $A\sim 130-195$ based on the pseudo DRHBc and RCHB mass tables denoted as red squares and blue squares, respectively. Stable nuclei (black square), experimentally known nuclei [8] (gray square), drip lines predicted by the pseudo DRHBc (red line) and RCHB (blue line) mass tables are also displayed.
@@
Figure 6 shows the differences of nuclear binding energy predicted by the pseudo DRHBc and RCHB mass tables. The differences of nuclear binding energy predictions are relatively small around neutron and proton magic numbers, while they become larger in the middle of the shell. In the simulation, most $r$-process path nuclei with $A=134-181$ are located in the $N=82-126$ region, where the pseudo DRHBc mass table predicts higher binding energies than the RCHB mass table due to the deformation effect. This leads to the differences of $r$-process abundances between these two mass models. The mass differences are mainly caused by the deformation effects. Fig. 7 shows the quadrupole deformation $\beta_{2}$ of even-$Z$ nuclei predicted by the DRHBc theory. The spherical and near-spherical nuclei are often found around neutron and proton magic numbers as expected, and the well-deformed nuclei often appear in the middle of the shell. Meanwhile the prolate-oblate shape changes can be found around $N=120$ for $50\leq Z\leq 82$. Notably, the $r$-process path with $A=134-181$ passes through the region for $N=82-126$. This indicates that the prediction of $r$-process path with $A=134-181$ is significantly influenced by the deformation effect.
@@
Figure 6: The binding energy differences for bound nuclei with $82\leq N\leq 126$ between the pseudo DRHBc calculations and RCHB calculations scaled by colors.
@@
As is evident from Eq. (10), the single neutron separation energy $S_{n}$ manifests in the exponent, rendering it the primary quantity that has a direct impact on the $r$-process simulations. Fig. 8 (a) shows the predicted neutron separation energies for the xenon isotopic chain based on the pseudo DRHBc and RCHB mass tables as functions of $A$. Note that in the present $r$-process simulations, the neutron flow with neutron number density $n_{n}=10^{24.5}$ cm-3 contributes the most to the abundance near $A=160$. Therefore, taking the neutron flow with $n_{n}=10^{24.5}$ cm-3 and the xenon ($Z=54$) isotopic chain as examples, we analyze the impact of the deformation effect on $r$-process nuclei around $A=160$. It is found that for the xenon isotopic chain, the neutron separation energies predicted by the pseudo DRHBc mass table decrease more rapidly with $A$. According to Eq. (10), as shown in Fig 8 (b), the abundance ratios of adjacent even-even nuclei predicted by the pseudo DRHBc mass table decrease more rapidly with $A$. Additionally, as the mass number $A$ increases, the abundance ratios of adjacent even-even nuclei decrease from several orders of magnitude above 1 to below 1. This indicates that as the neutron number increases, the abundance of even-even nuclei gradually increases until reaching a maximum and then decreases. Fig. 8 (c) shows the relative abundance distribution of even-even nuclei on the xenon isotopic chain. It is found that, compared to the results of the RCHB mass table, the maximum of the abundance distribution, which corresponds to the $r$-process path nuclei, based on the pseudo DRHBc mass table is located at smaller mass numbers. This is the reason that the $r$-process abundances based on pseudo DRHBc mass table are higher for $A=148-165$ compared to those based on the RCHB mass table, as shown in Fig. 4.
Figure 8: (a) The neutron separation energies of the xenon isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number $A$. (b) The abundance ratio of adjacent even-even nuclei along the xenon isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number $A$, under astrophysical conditions with neutron number density $n_{n}=$ $10^{24.5}\mathrm{~{}cm}^{-3}$ and temperature $T\gtrsim 1.5\mathrm{~{}K}$. (c) Relative abundance distribution of even-even nuclei along the xenon isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number $A$, before the neutrons freeze-out.
As for the $r$-process nuclei around $A=175$, the neutron flow with neutron number density $n_{n}=10^{24}$ cm-3 and the neodymium ($Z=60$) isotopic chain are taken to analyze the impact of deformation effect. Figure 9 (b) represents the predicted neutron separation energies for the neodymium isotopic chain based on the pseudo DRHBc and RCHB mass tables as functions of mass number $A$. It is found that the pseudo DRHBc mass table with rotational corrections predicts a sudden decrease in neutron separation energy at $A=180$. This sudden change arises because the DRHBc calculation neglects rotational corrections for nuclei with $|\beta_{2}|<0.05$ to avoid numerical challenges [27, 77]. As shown in Fig. 9 (a), the quadrupole deformations $\beta_{2}$ from the DRHBc calculations for neodymium isotopes drop below 0.05 at $A=180$, causing the sudden reduction in neutron separation energy for this isotope. Consequently, this sharp decline reduces the relative abundance distribution at $A=180$. Following the neutron freeze-out stage, this phenomenon influences the $r$-process abundance pattern within the $A=170-178$ region, primarily as a result of the $\beta$-decay processes that are accompanied by neutron emission. In addition, it is found that the neutron separation energies predicted by the pseudo DRHBc mass table show a non-monotonic behavior as $A$ increases and represent a bump around $A=175-182$. This leads to abundance ratios of adjacent even-even isotopes, as shown in Fig. 9(c), being less than 1 around $A=172-176$ and larger than 1 for $A>176$, thus forming a minimum near $A=176$. Correspondingly, the relative abundances predicted by the pseudo DRHBc mass table exhibit a trough around $A=176$, as shown in Fig. 9 (d). This simulated abundance trough, due to non-monotonic behavior of neutron separation energies, is also noted in Ref. [102]. The rapid shape transitions from a prolate configuration to an oblate one may explain this non-monotonic behavior. As shown in Fig. 9 (a), the quadrupole deformation $\beta_{2}$ for the neodymium isotopes shows a rapid shape transitions from prolate to oblate at $A=175$, corresponding to the predicted non-monotonic neutron separation energy changes.
@@
Figure 9: (a) The nuclear quadrupole deformations $\beta_{2}$ of neodymium isotopes predicted by the DRHBc theory as a function of mass number $A$. (b) The neutron separation energies of the neodymium isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number $A$. (c) The abundance ratio of adjacent even-even nuclei along the neodymium isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number $A$, under astrophysical conditions with neutron number density $n_{n}=$ $10^{24}\mathrm{~{}cm}^{-3}$ and temperature $T=1.5\mathrm{~{}K}$. (d) Relative abundance distribution of even-even nuclei along the neodymium isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number $A$.
@@
The abundance trough around $A=170$ from classical $r$-process simulation with the pseudo DRHBc mass table is significant. As it is very different from the observed solar $r$-process abundances, here we provide two more remarks:
@@
    First, from the view of nuclear structure, the rapid shape transitions have led to the abundance trough around $A=170$. Rapid shape transitions occur due to the shape-coexistence of prolate and oblate configurations for nuclei in this region. In this sense, incorporating the triaxial deformation effects or the shape-mixed beyond-mean-field effects could potentially resolve this abundance trough.
@@
    Third, from the view of methodology, this abundance trough might only appear in the present pseudo DRHBc mass table. Although the pseudo DRHBc mass predictions are very close to the real DRHBc results, the method of estimation employed in the present work is still not expected to be as effective as the real DRHBc calculations in dealing with the odd-even effects, which could lead to certain deviations in the results. One can expect the release of the full DRHBc mass table and test whether the abundance trough would appear.
V Summary
@@
In summary, based on the available even-$Z$ part of the DRHBc mass table, the masses of odd-$Z$ nuclei are estimated by approximating the odd-even mass differences with average pairing gaps. A pseudo DRHBc mass table is constructed, and applied to $r$-process simulation.
@@
The estimation is realized by expressing the mass of an odd nucleus as a function of its neighboring even ones’ masses and odd-even mass differences, where the latter are approximated by the average pairing gaps. This treatment is examined by taking all even-odd nuclei with $8\leqslant Z\leqslant 120$ as examples and estimating their binding energies based on the DRHBc results of even-even nuclei. It is found that the estimated binding energies of even-odd nuclei can effectively reproduce the self-consistent DRHBc calculation results. Then we estimate the binding energies of odd-$Z$ nuclei based on the DRHBc results of even-$Z$ nuclei, and combine them to construct a pseudo DRHBc mass table for all the nuclei with $8\leqslant Z\leqslant 120$. The rms deviation of binding energy from available data $\sigma=2.52$ MeV (without rotational correction) and 1.50 MeV (with rotational correction). The one/two-neutron/proton separation energies as well as their accuracies are also calculated and found to be close to the real DRHBc results from self-consistent calculations.
@@
The $r$-process simulation is performed based on the obtained pseudo DRHBc mass table, and the impact of nuclear deformation effect is analyzed via comparing with results from the RCHB mass table. The simulations from both mass tables predict abundance peaks near $A=130,195$ and an abundance trough around $A=170$. Compared to the RCHB mass table, the pseudo DRHBc mass table predicts higher abundances in $A=148-165$ and lower abundances in $A=170-180$. These differences can be explained by the $r$-process path nuclei predicted by the pseudo DRHBc mass tables; namely, around $A=160$, they are closer to the stability line, while around $A=175$ there is a gap. It is found that the differences in the $r$-process path nuclei around $A=160$ are caused by the faster decrease of neutron separation energies predicted by the pseudo DRHBc mass table along mass number $A$ on the $Z=54-58$ isotope chains. The absence of $r$-process path nuclei around $A=175$ is attributed to the non-monotonic behavior of neutron separation energy and also influenced by the treatment of rotational corrections in the present DRHBc calculations. Detailed analysis indicates that both of these two differences originate from the deformation effects.
@@
For perspective, full-network dynamical $r$-process simulations based on the upcoming full DRHBc mass table will be essential to clarify the abundance trough around $A=170$. Incorporating the triaxial deformation and the beyond-mean-field effects in the DRHBc calculations for several key nuclei are also crucial, as more sufficient descriptions for nuclei with rapid shape transitions could be provided. Further improvements are expected based on these extensions.
Acknowledgements.

### v2

@@ 4 $R$-process simulation
The obtained pseudo DRHBc mass table in Section 3 is applied in the $r$-process simulation, to study the impact of deformation effects in the $r$-process. The $r$-process simulation is performed using the classical $r$-process model, where nuclear masses are taken from the pseudo DRHBc mass table, if the experimental data Wang2021CPC are not available. For the $\beta$-decay rates, the empirical formula Zhou2017NPR using decay energies from the pseudo DRHBc calculations is employed together with experimental data Audi2003NPA . The astrophysical trajectory, i.e., the weight $\omega$ and the irradiation time $\tau$ of neutron flows, is determined by fitting the obtained abundances to the solar $r$-process abundances Sneden2008ARNPS at a temperature of T=1.5GKT=1.5\ $\mathrm${GK}. For comparison, corresponding $r$-process simulation based on the RCHB mass table Xia2018ADNDT is also carried out.
@@
Figure 6: The $r$-process abundances from classical $r$-process simulations based on pseudo DRHBc and RCHB mass tables as functions of the mass number AA. The solar $r$-process abundances Sneden2008ARNPS are displayed with black dots.
The $r$-process abundances with pseudo DRHBc and RCHB mass tables are shown in Fig. 6. It is found that the simulated abundances with both the RCHB and pseudo DRHBc mass tables exhibit abundance peaks near $A=130,195$ and an abundance trough around $A=170$. The $r$-process abundances based on pseudo DRHBc mass table are higher for A=148165A=148-165 and lower for A=170178A=170-178 compared to those based on the RCHB mass table. This indicates the impact from the deformation effects. We note that discrepancies exist between Solar $r$-process abundances and the ones based on pseudo DRHBc or RCHB mass tables. The reasons for this can be multifaceted. For example, the triaxial and octupole degrees of freedom, as well as beyond-mean-field effects, are not considered in current treatments, which may lead to deviations in nuclear masses. Beyond the approximations in symmetry, uncertainties also exist in density functional calculations due to different fitting protocols Agbemava2014PRC . The approximations in the classical $r$-process simulations, e.g., waiting point approximation and sudden freeze-out approximation, can also contribute to such discrepancies.
Figure 7: The $r$-process abundances from classical $r$-process simulations based on the masses of even-$Z$ nuclei from the DRHBc mass table and the masses of odd-$Z$ nuclei from estimations using different pairing gaps, i.e., the microscopic average pairing gaps (pseudo DRHBc), the empirical gaps $\Delta=12A^{-1/2}$, and $\Delta=34A^{-3/4}$. The solar $r$-process abundances Sneden2008ARNPS are displayed with black dots.
Before proceeding to explore the deformation effect, considering that the input masses of odd-$Z$ nuclei are estimated based on the average pairing gaps in Section 3.3, it would be important to examine whether $r$-process simulation is sensitive to the choice of pairing gaps. In Fig. 7, the $r$-process abundances are shown, with the masses of odd-$Z$ nuclei from estimations using the microscopic average pairing gaps (pseudo DRHBc) and the empirical gaps. It is found that the $r$-process abundances from empirical gaps exhibit qualitatively the same behaviors as the pseudo DRHBc results. This indicates that these pairing-gap choices are sufficient to support the estimation of masses for nuclei that have not been explicitly calculated.
@@
The most important nuclei for the $r$-process are located in the $r$-process path. The $r$-process paths for nuclei with $A\sim 130-195$ based on the pseudo DRHBc and RCHB mass tables are displayed in Fig. 8. It is found that the $r$-process path nuclei around $A=160$ predicted by the pseudo DRHBc mass table are closer to the stability line than the ones predicted by the RCHB mass table. Since the beta decay half-lives are generally shorter when the nuclei are moving toward the neutron-rich direction, the decay half-lives of $r$-process path nuclei around $A=160$ based on the pseudo DRHBc mass table are longer than those based on the RCHB mass table. This leads to increased accumulation of abundances on the path nuclei. It is therefore understandable that the $r$-process abundances based on the pseudo DRHBc mass table are larger for A=148165A=148-165 than the ones based on the RCHB mass table. On the other hand, it is noted that the $r$-process path predicted by the pseudo DRHBc mass table has a gap around $A=175$, which leads to a deficiency in the simulated abundances in this region.
@@
Figure 8: The $r$-process path for nuclei with $A\sim 130-195$ based on the pseudo DRHBc and RCHB mass tables denoted as red filled squares and blue filled circles, respectively. Stable nuclei (black filled square), experimentally known nuclei Wang2021CPC (gray filled square), drip lines predicted by the pseudo DRHBc (red line) and RCHB (blue line) mass tables are also displayed.
@@
Figure 9 shows the differences of nuclear binding energy predicted by the pseudo DRHBc and RCHB mass tables. The differences of nuclear binding energy predictions are relatively small around neutron and proton magic numbers, while they become larger in the middle of the shell. In the simulation, most $r$-process path nuclei with A=134181A=134-181 are located in the N=82126N=82-126 region, where the pseudo DRHBc mass table predicts higher binding energies than the RCHB mass table due to the deformation effect. This leads to the differences of $r$-process abundances between these two mass models. The mass differences are mainly caused by the deformation effects. Figure 10 shows the quadrupole deformation $\beta_{2}$ of even-$Z$ nuclei predicted by the DRHBc theory. The spherical and near-spherical nuclei are often found around neutron and proton magic numbers as expected, and the well-deformed nuclei often appear in the middle of the shell. Meanwhile the prolate-oblate shape changes can be found around $N=120$ for $50\leq Z\leq 82$. Notably, the $r$-process path with A=134181A=134-181 passes through the region for N=82126N=82-126. This indicates that the prediction of $r$-process path with A=134181A=134-181 is significantly influenced by the deformation effect.
@@
Figure 9: The binding energy differences for bound nuclei with $82\leq N\leq 126$ between the pseudo DRHBc calculations and RCHB calculations scaled by colors. The $r$-process path for nuclei with $A\sim 130-195$ based on the pseudo DRHBc and RCHB mass tables denoted as red filled squares and blue filled circles, respectively.
@@
As is evident from Eq. (11), the single neutron separation energy $S_{n}$ manifests in the exponent, rendering it the primary quantity that has a direct impact on the $r$-process simulations. Taking the neutron flow with $n_{n}=10^{24.5}$ cm-3 and the xenon ($Z=54$) isotopic chain as examples, we analyze the deformation effect on the $S_{n}$ of $r$-process nuclei around $A=160$. The xenon isotopes show prolate shapes at A=140159A=140-159 and oblate shapes for A=160170A=160-170, as shown in Fig. 11(a). Figure 11(b) shows the predicted neutron separation energies for the xenon isotopic chain based on the pseudo DRHBc and RCHB mass tables as functions of $A$. After including the deformation effect, the neutron separation energies predicted by the pseudo DRHBc mass table decrease more rapidly with $A$. Note that in the present $r$-process simulations, the neutron flow with neutron number density $n_{n}=10^{24.5}$ cm-3 contributes the most to the abundance near $A=160$. According to Eq. (11), as shown in Fig. 11(c), the abundance ratios of adjacent even-even nuclei predicted by the pseudo DRHBc mass table decrease more rapidly with $A$. Additionally, as the mass number $A$ increases, the abundance ratios of adjacent even-even nuclei decrease from several orders of magnitude above 1 to below 1. This indicates that as the neutron number increases, the abundance of even-even nuclei gradually increases until reaching a maximum and then decreases. Figure 11(d) shows the relative abundance distribution of even-even nuclei on the xenon isotopic chain. It is found that, compared to the results of the RCHB mass table, the maximum of the abundance distribution, which corresponds to the $r$-process path nuclei, based on the pseudo DRHBc mass table is located at smaller mass numbers. This is the reason that the $r$-process abundances based on pseudo DRHBc mass table are higher for A=148165A=148-165 compared to those based on the RCHB mass table, as shown in Fig. 6.
@@
Figure 11: (a) The nuclear quadrupole deformations $\beta_{2}$ of xenon isotopes predicted by the DRHBc theory as a function of mass number AA. (b) The neutron separation energies of the xenon isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number AA. (c) The abundance ratio of adjacent even-even nuclei along the xenon isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number AA, under astrophysical conditions with neutron number density $n_{n}=$ $10^{24.5}\mathrm{~cm}^{-3}$ and temperature $T\gtrsim 1.5\mathrm{~K}$. (d) Relative abundance distribution of even-even nuclei along the xenon isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number AA, before the neutrons freeze-out.
As for the $r$-process nuclei around $A=175$, the neutron flow with neutron number density $n_{n}=10^{24}$ cm-3 and the neodymium ($Z=60$) isotopic chain are taken to analyze the impact of deformation effect. Figures 12(a) and (b) represent the predicted quadrupole deformations $\beta_{2}$ and neutron separation energies for the neodymium isotopic chain, respectively. It is found that the pseudo DRHBc mass table with rotational corrections predicts a sudden decrease in neutron separation energy at $A=180$. This sudden change arises because the DRHBc calculation neglects rotational corrections for nuclei with $|\beta_{2}|<0.05$ since the cranking approximation used is not suitable for near-spherical nuclei Zhang2022ADNDT ; Guo2024ADNDT , as has been discussed in Section 2. As shown in Fig. 12(a), the quadrupole deformations $\beta_{2}$ from the DRHBc calculations for neodymium isotopes drop below 0.05 at $A=180$, causing the sudden reduction in neutron separation energy for this isotope. Consequently, this sharp decline reduces the relative abundance distribution at $A=180$. Following the neutron freeze-out stage, this phenomenon influences the $r$-process abundance pattern within the A=170178A=170-178 region, primarily as a result of the $\beta$-decay processes that are accompanied by neutron emission. In addition, it is found that the neutron separation energies predicted by the pseudo DRHBc mass table show a non-monotonic behavior as $A$ increases and represent a bump around A=175182A=175-182. This leads to abundance ratios of adjacent even-even isotopes, as shown in Fig. 12(c), being less than 1 around A=172176A=172-176 and larger than 1 for $A>176$, thus forming a minimum near $A=176$. Correspondingly, the relative abundances predicted by the pseudo DRHBc mass table exhibit a trough around $A=176$, as shown in Fig. 12(d). It is noted in Fig. 11(a) that Xe isotopes also show $|\beta_{2}|$ crossing $0.05$ near $A=180$, which leads to a trough in $S_{2n}/2$ and Y(A)/Y(A2)Y(A)/Y(A-2), but since their Y(A)/Y(A2)Y(A)/Y(A-2) values are significantly smaller than $1$ and $A=180$ reaches the neutron drip line for Xe, the abundances of such extremely neutron-rich isotopes are very low, and therefore, the trough in abundance for Xe isotopes does not show up. Similar trough has been noted in Ref. Martin2016PRL , where the abundance based on SLy4 mass model shows a trough around $A=180$, which does not appear in the results based on SkM* and UNEDF0 mass models. This trough was also attributed to the nonmonotonic behavior of $S_{2n}$, which was traced back to rapid shape transitions from prolate to oblate. As shown in Fig. 12 (a), the quadrupole deformation $\beta_{2}$ for the neodymium isotopes shows a rapid shape transitions from prolate to oblate at $A=175$, corresponding to the predicted non-monotonic neutron separation energy changes.
Figure 12: (a) The nuclear quadrupole deformations $\beta_{2}$ of neodymium isotopes predicted by the DRHBc theory as a function of mass number AA. (b) The neutron separation energies of the neodymium isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number AA. (c) The abundance ratio of adjacent even-even nuclei along the neodymium isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number AA, under astrophysical conditions with neutron number density $n_{n}=$ $10^{24}\mathrm{~cm}^{-3}$ and temperature $T=1.5\mathrm{~K}$. (d) Relative abundance distribution of even-even nuclei along the neodymium isotope chain predicted by the pseudo DRHBc and RCHB mass tables as functions of mass number AA.
@@
The abundance trough around $A=170$ from classical $r$-process simulation with the pseudo DRHBc mass table is significant. As it is very different from the observed solar $r$-process abundances, here we provide three more remarks:
@@
    First, from the view of nuclear structure, the rapid shape transitions have led to the abundance trough around $A=170$. Rapid shape transitions occur due to the shape-coexistence of prolate and oblate configurations for nuclei in this region. Such shape coexistence might correspond to the triaxial deformation Zhang2023PRC_TRHBc ; Zhang2025arXiv . Incorporating the triaxiality is expected to smooth the evolution of nuclear shapes, leading to consistently smoother evolutions of $S_{2n}$ and $Y(A)$ and resolving the abundance trough. In addition, the shape-mixed beyond-mean-field effects Sun2022CPC could also potentially improve the abundance description at this trough.
@@
    Third, from the view of methodology, it has been concluded above that the $r$-process abundances are not very sensitive to the details of pairing gaps. Therefore, it is reasonably inferred that this trough might still exist in the abundance based on the full DRHBc mass table in the future, due to the shape transitions. One can expect the examination for this trough after the release of the full DRHBc mass table.
5 Summary
@@
In summary, based on the available even-$Z$ part of the DRHBc mass table with the density functional PC-PK1 and the density-dependent zero-range pairing force, the masses of odd-$Z$ nuclei are estimated by approximating the odd-even mass differences with microscopic average pairing gaps. A pseudo DRHBc mass table is constructed, and applied to $r$-process simulation.
@@
The estimation is realized by expressing the mass of an odd nucleus as a function of its neighboring even ones’ masses and odd-even mass differences, where the latter are approximated by the microscopic average pairing gaps. This treatment is examined by taking all even-odd nuclei with $8\leqslant Z\leqslant 120$ as examples and estimating their binding energies based on the DRHBc results of even-even nuclei. It is found that the estimated binding energies of even-odd nuclei can effectively reproduce the self-consistent DRHBc calculation results. The comparisons with the results based on the empirical pairing gaps $12A^{-1/2}$ and $34A^{-3/4}$ show that the global behavior of the binding energy is not very sensitive to the details in the evolution of pairing gaps. Then we estimate the binding energies of odd-$Z$ nuclei based on the DRHBc results of even-$Z$ nuclei, and combine them to construct a pseudo DRHBc mass table for all the nuclei with $8\leqslant Z\leqslant 120$. The rms deviation of binding energy from available data $\sigma=2.55$ MeV (without rotational correction) and 1.47 MeV (with rotational correction). The one/two-neutron/proton separation energies as well as their accuracies are also calculated and found to be close to the real DRHBc results from self-consistent calculations.
@@
The $r$-process simulation is performed based on the obtained pseudo DRHBc mass table. The pairing gaps chosen are sufficient to support the estimation of masses for nuclei that have not been explicitly calculated. The impact of nuclear deformation effect is analyzed via comparing with results from the RCHB mass table. The simulations from both mass tables predict abundance peaks near $A=130,195$ and an abundance trough around $A=170$. Compared to the RCHB mass table, the pseudo DRHBc mass table predicts higher abundances in A=148165A=148-165 and lower abundances in A=170180A=170-180. These differences can be explained by the $r$-process path nuclei predicted by the pseudo DRHBc mass tables; namely, around $A=160$, they are closer to the stability line, while around $A=175$ there is a gap. It is found that the differences in the $r$-process path nuclei around $A=160$ are caused by the faster decrease of neutron separation energies predicted by the pseudo DRHBc mass table along mass number $A$ on the Z=5458Z=54-58 isotope chains. The absence of $r$-process path nuclei around $A=175$ is attributed to the non-monotonic behavior of neutron separation energy and also influenced by the treatment of rotational corrections in the present DRHBc calculations. Detailed analysis indicates that both of these two differences originate from the deformation effects.
@@
For perspective, full-network dynamical $r$-process simulations based on the upcoming full DRHBc mass table will be essential to clarify the abundance trough around $A=170$. Incorporating the triaxial deformation and the beyond-mean-field effects in the DRHBc calculations for several key nuclei are also crucial, as more sufficient descriptions for nuclei with rapid shape transitions could be provided. Further improvements are expected based on these extensions. In addition, a more detailed comparison with other microscopic or semi-empirical mass models in terms of $r$-process abundance results is to be conducted in future investigations.
Acknowledgments